Will you please show detail of how this problem is solved and the correct answer? Thank you
A total of $15,000.00 was invested in two accounts, one at 4% and the other at 3% simple intrest. How much was invested in each account if the total earned was $550.00 per year?

Answer by brucewill(101)

Let x = the amount of 3% money
Then (15000-x) = the amount fo 4% money.
Interest = Principal * rate * time.
0.03x + 0.04x(15000-x) = 550
---
0.03x + 600 - 0.04x = 550
---
-0.01x = -50
===
x = 5000 the amount of 3% money.
Therefore, 15000 - 5000 = 10000 the amount of 4% money.
Proof:
5000 * 3% = 150
10000 * 4% = 400, sums to 550.